# How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library

### v3

@@ Abstract
With the rise of Large Language Models (LLMs) in recent years, abundant new opportunities are emerging, but also new challenges, among which contamination is quickly becoming critical. Business applications and fundraising in AI have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a major issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes the entire progress in the field of NLP, yet, there remains a lack of methods on how to efficiently detect contamination. In this paper, we survey all recent work on contamination detection with LLMs, and help the community track contamination levels of LLMs by releasing an open-source Python library named LLMSanitize implementing major contamination detection algorithms, available here: https://github.com/ntunlp/LLMSanitize
How Much are Large Language Models Contaminated? A Comprehensive Survey and the LLMSanitize Library
Mathieu Ravaut1,2, Bosheng Ding1, Fangkai Jiao1,2, Hailin Chen1,3, Xingxuan Li1, Ruochen Zhao1, Chengwei Qin1, Caiming Xiong3, Shafiq Joty1,3† 1 Nanyang Technological University, Singapore 2 Institute of Infocomm Research (I2R), ASTAR, Singapore 3 Salesforce Research {mathieuj001,bosheng001,fangkai002,hailin001,xingxuan001,ruochen002,chengwei003}@e.ntu.edu.sg {cxiong,sjoty}@salesforce.com
**footnotetext: Authors contributed equallyfootnotetext: Work done full-time at Salesforce
@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), Large Language Models (LLMs) have emerged as pivotal tools, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023a; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored output (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness and reliability.
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points. Additionally, we highlight the following orthogonal dimension to the duality explained above: if we consider whether the contaminated data contains only inputs or both inputs and labels, contamination includes input contamination and input+label contamination. A third dimension to the problem is the level of access to the LLM itself in closed-data setups: white-box detection where we have full access to the model, e.g., a local model, and black-box detection where only the API access is available, e.g., GPT-4 (Achiam et al., 2023).
@@
Matching, fill=orange!30, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), PaLM (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023), Dodge et al. (2021); Li (2023b); Deng et al. (2023), models ] ] [Embeddings
Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), Riddell et al. (2024), models ] ] [Model-based Open-
Data Contamination Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023b), models] ] ] [Closed-Data
Contamination Detection, fill=yellow!30, domain_box [Performance
Analysis, fill=orange!30, domain_box [Roberts et al. (2023); Li and Flanigan (2023); Jiang et al. (2024b), models ] ] [Model
Memorization, fill=orange!30, domain_box [TS-Guessing (Deng et al., 2023), Elangovan et al. (2021); Magar and Schwartz (2022); Gemma Team Google DeepMind (2024); Nasr et al. (2023); Weller et al. (2023); Chang et al. (2023); Ranaldi et al. (2024), models ] ] [Model
Confidence, fill=orange!30, domain_box [Min-KProb (Shi et al., 2024), CDD and TED (Dong et al., 2024), Li (2023a); Oren et al. (2023); Xu et al. (2024b), models ] ] [Model-based Closed-
Data Contamination Detection, fill=orange!30, domain_box [Guided prompting (Golchin and Surdeanu, 2023b), Tu et al. (2024), models] ] ] ]
Figure 1: Classification of the different types of contamination detection methods.
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In high-stakes scenarios where imprecision can have dramatic consequences, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination.
This landscape necessitates a comprehensive survey and a shared integrated resource of contamination detection in LLMs. We delineate the scope and nature of contamination detection challenges, identifying their types. In doing so, we not only contribute to the academic and practical understanding of data contamination issues but also highlight the pressing need for strategies that mitigate these risks.
@@
  • To the best of our knowledge, we are the first to thoroughly review the field of contamination detection in LLMs.
  • We categorize contamination detection into two broad types of use cases vastly differing in the techniques involved, namely open-data and closed-data contamination detection, and review all existing work in each category. We refer to Figure 1 for an overview of our proposed classification of contamination detection methods.
  • We introduce LLMSanitize, an open-source Python library aimed at helping the community centralize a shared implementation of all contamination detection methods.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to open-data contamination detection. Section 3 addresses the issue of closed-data contamination detection, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices. Section 5 briefly introduces our library LLMSanitize implementing contamination detection algorithms, which is further described in Appendix.
@@ 2.1 String Matching
Several LLM pre-training reports assess contamination levels by using string matching techniques. GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. An average of 3.2% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training is discovered. GPT-3 (Brown et al., 2020) scans data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). Some Wikipedia language modeling benchmarks, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018) are nearly entirely contaminated. PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Both GPT-3 and PaLM authors conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees. GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data by computing 50-characters substring collisions. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results. Llama-2 (Touvron et al., 2023) refines the methods above by using tokens instead of words to measure contamination, and asserts that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. From there, the contamination level of a sample is defined as its percentage of contaminated tokens.
@@
Other studies analyze overlap between training sets and evaluation sets to assess contamination levels. Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the C4 pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%. Li (2023b) compute the METEOR (Banerjee and Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and consider those with scores over 0.75 as contaminated. Contamination levels ranging from 1% to 47% across popular benchmarks are identified. Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then splitting retrieved documents into 13-gram chunks, and computing overlap metrics between these chunks and the evaluation data point’s chunks. TruthfulQA (Lin et al., 2021) exhibits high overlap with the pre-training datasets.
@@ 2.2 Embeddings Similarity
Lee et al. (2023) prevent contamination in their Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers and Gurevych, 2019)) greater than 80% against any training item. Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their CodeExercises dataset and the evaluation set HumanEval (Chen et al., 2021). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from (Riddell et al., 2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). Widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 Model-based Open-Data Contamination Detection
Embeddings similarity is limited by its need to choose an appropriate similarity threshold. To improve robustness, some methods attempt to use LLMs themselves to detect contamination in LLMs, leveraging their high performance in paraphrasing and paraphrase detection (Witteveen and Andrews, 2019; Abaskohi et al., 2023).
Yang et al. (2023b) argue that contamination checks become challenging because of the presence of “rephrased samples”, which have the same semantics but different surface form as the original sample. Therefore, they first use embeddings similarity search to get the top-k similar samples with a given test sample, then prompt a strong LLM (namely, GPT-4) to examine whether any of the top-k samples is too close to the test case. Results show that this LLM Decontaminator algorithm works significantly better than existing methods at tagging contaminated samples.

### v4

# A Comprehensive Survey of Contamination Detection Methods in Large Language Models

@@ Abstract
With the rise of Large Language Models (LLMs) in recent years, abundant new opportunities are emerging, but also new challenges, among which contamination is quickly becoming critical. Business applications and fundraising in Artificial Intelligence (AI) have reached a scale at which a few percentage points gained on popular question-answering benchmarks could translate into dozens of millions of dollars, placing high pressure on model integrity. At the same time, it is becoming harder and harder to keep track of the data that LLMs have seen; if not impossible with closed-source models like GPT-4 and Claude-3 not divulging any information on the training set. As a result, contamination becomes a major issue: LLMs’ performance may not be reliable anymore, as the high performance may be at least partly due to their previous exposure to the data. This limitation jeopardizes real capability improvement in the field of NLP, yet, there remains a lack of methods on how to efficiently detect contamination. In this paper, we survey all recent work on contamination detection with LLMs, analyzing their methodologies and use cases to shed light on the appropriate usage of contamination detection methods. Our work calls the NLP research community’s attention into systematically taking into account contamination bias in LLM evaluation.
@@ 1 Introduction
In the rapidly evolving landscape of artificial intelligence (AI), large language models (LLMs) have emerged as a pivotal tool, driving innovation across a wide spectrum of applications, from natural language processing (NLP) and automated content creation (Achiam et al., 2023; Betker et al., 2023) to complex decision-making systems and autonomous agents (Wei et al., 2022; Li et al., 2023a; Yang et al., 2023b; Wu et al., 2023). At their core, these models rely on extensive datasets (Raffel et al., 2020; Gao et al., 2020) to learn about language and the world, and generate responses that are increasingly indistinguishable from human-authored writing (Ouyang et al., 2022a). However, the integrity of these datasets is paramount, as any contamination can significantly impair the models’ effectiveness, reliability, and generalization. In this study, we define the phenomenon of contamination as any leakage of the evaluation data within the training data.
Contamination poses a multifaceted challenge, threatening not only the technical accuracy of LLMs but also their ethical and commercial viability. In high-stakes scenarios where imprecision can have dramatic consequences, such as medical diagnosis, legal advice, or financial services, the repercussions of relying on contaminated data can be profound. Moreover, the allure of leveraging LLM outputs to attract investment underscores a pressing commercial dimension. As businesses increasingly integrate AI-driven insights into their strategic planning and operational decisions, the assurance of data purity becomes intertwined with potential market success and valuation. Lastly, in the current intense race to build the most powerful LLM (Touvron et al., 2023b; Team et al., 2023; Jiang et al., 2024a; Bai et al., 2023; Bi et al., 2024; Young et al., 2024; Gemma Team Google DeepMind, 2024; Chen et al., 2023), the community is struggling to settle on a fixed subset of benchmarks as LLMs performance increases fast (Clark et al., 2018; Zellers et al., 2019; Hendrycks et al., 2020; Lin et al., 2021; Sakaguchi et al., 2021; Cobbe et al., 2021), an issue which is further fueled by underlying contamination. This landscape necessitates a comprehensive survey of contamination detection in LLMs.
Detecting contamination in a LLM is a broad issue that we divide into open-data contamination detection and closed-data contamination detection depending on the nature of the training set. On the one hand, open-data refers to the scenario where the LLM pre-training data is known, enabling direct comparisons with the evaluation dataset. On the other hand, closed-data contamination refers to the more and more prevalent use case of an unknown pre-training set. This latter case is more challenging and necessitates an examination of the LLM’s behavior on evaluation data points. Contamination may span only part of a (x, y) data point where x is the input and y the corresponding label. If we consider whether the contaminated data contains only x or both x and y, contamination can be divided into input contamination and input+label contamination. Detecting contamination is further complicated by varying level of access to the LLM. With white-box access, the LLM and all weights are available, e.g., a local model. In gray-box access, model weights are not available to the researcher but output token probability distributions are. Most closed-source setups, e.g., GPT-4, follow a black-box setup where only the output text can be accessed from the API. We will review in this study contamination detection methods specifically tailored for each of these levels of model access.
@@
Matching, fill=orange!30, domain_box [Instance-Level
String Matching, fill=red!20, domain_box [GPT-2 (Radford et al., 2019), GPT-3 (Brown et al., 2020), PaLM (Anil et al., 2023), GPT-4 (Achiam et al., 2023), Llama-2 (Touvron et al., 2023b), Llama-3 (Dubey et al., 2024), Qwen-2.5-Coder (Hui et al., 2024), Longest-Match (Singh et al., 2024), models ] ] [Dataset-Level
String Matching, fill=red!20, domain_box [Dodge et al. (2021); Li (2023b); Deng et al. (2023), models ] ] ] [Embeddings
Similarity, fill=orange!30, domain_box [Platypus (Lee et al., 2023), Phi-1 (Gunasekar et al., 2023), Riddell et al. (2024), models ] ] [Paraphrase
Detection, fill=orange!30, domain_box [LLM Decontaminator (Yang et al., 2023c), models] ] ] [Closed-Data
Contamination Detection, fill=yellow!30, domain_box [Dataset-level, fill=orange!30, domain_box [Performance
Analysis, fill=red!30, domain_box [Train-Test Performance
Gap, fill=purple!20, domain_box [CAP Zhao et al. (2024), models ] ] [Controlled Training
Setup, fill=purple!20, domain_box [Jiang et al. (2024b); Palavalli et al. (2024), models ] ] [Performance Change Through Time, fill=purple!20, domain_box [Aiyappa et al. (2023); Roberts et al. (2023); Li & Flanigan (2023); Cao et al. (2024), models ] ] ] ] [Instance-level, fill=orange!30, domain_box [Membership-Inference
Attacks (MIA), fill=red!30, domain_box [Reference-based
MIA, fill=purple!20, domain_box [RECALL (Xie et al., 2024), MaxNorm(Meeus et al., 2024a), MIA-Tuner (Fu et al., 2024a), PREMIA (Feng et al., 2024), models ] ] [Reference-free
MIA, fill=purple!20, domain_box [SPV-MIA (Fu et al., 2023), CAMIA (Chang et al., 2024), S2-MIA (Li et al., 2024), EM-MIA (Kim et al., 2024), models ] ] ] [Model
Memorization, fill=red!30, domain_box [Quantifying
Memorization, fill=purple!20, domain_box [Elangovan et al. (2021); Magar & Schwartz (2022); Gemma Team Google DeepMind (2024), models ] ] [Prompt
Engineering, fill=purple!20, domain_box [Nasr et al. (2023), According To prompting (Weller et al., 2023), Guided prompting (Golchin & Surdeanu, 2023a), Local order quiz (Samuel et al., 2024), models ] ] [Cloze
Tasks, fill=purple!20, domain_box [TS-Guessing (Deng et al., 2023), Chang et al. (2023); Ranaldi et al. (2024), models ] ] ] [Model
Confidence, fill=red!30, domain_box [Likelihood, fill=purple!20, domain_box [Min-KProb (Shi et al., 2024), CDD and TED (Dong et al., 2024), DC-PDD (Zhang et al., 2024b), Min-K%++ (Zhang et al., 2024a), SURP (Zhang & Wu, 2024), Li (2023a); Oren et al. (2023); Xu et al. (2024b), models ] ] [Probing Hidden Layers, fill=purple!20, domain_box [DICE (Tu et al., 2024), Liu et al. (2024), models ] ] ] ] ] ]
Figure 1: Classification of contamination detection methods reviewed in this paper.
@@
  • To the best of our knowledge, we are the first to thoroughly review the field of contamination detection in LLMs. We review all reported contamination detection methods up until early 2025, totaling more than 50 detection techniques and more than 100 relevant papers.
  • We categorize contamination detection into two broad types of use cases vastly differing in the techniques involved, namely open-data and closed-data contamination detection, and review all existing work in each category. Our taxonomy further classifies detection techniques into finer-grained categories, each following their own assumptions and model requirements.
  • Through our detailed classification and review of the types of contamination detection, we not only contribute to the academic and practical understanding of data contamination issues but we also highlight the pressing need for strategies that mitigate these risks.
We refer to Figure 1 for an overview of our proposed classification of contamination detection methods. The concurrent work of Palavalli et al. (2024) proposes a taxonomy of the different types of contamination. We review the detection of contamination at large, unrestricted to a specific form of contamination such as a particular noising then injection of test data points into the training set. The concurrent works of Deng et al. (2024) and Fu et al. (2024b) also survey contamination in LLMs. Compared to concurrent surveys, we dive deep with a unique focus on contamination detection methodologies and provide the most comprehensive overview of current papers, enabling enabling researchers to quickly grasp the existing research field of contamination detection in LLMs as of early 2025.
The rest of this paper is structured as follows: Section 2 explores methods and findings related to open-data contamination detection. Section 3 addresses the issue of closed-data contamination detection, where models may be evaluated on data that they have previously seen. Section 4 discusses the current and future challenges in the field, including best practices, and reviews all new datasets proposed to detect contamination.
@@ 2.1.1 Instance-Level String Matching
Several LLM pre-training reports assess contamination levels by using string matching techniques. GPT-2 (Radford et al., 2019) calculates contamination as the percentage of 8-grams from a particular evaluation set that are also found in the WebText training set. An average of 3.2% overlap between common LM datasets’ test sets and the WebText dataset used for pre-training is discovered, raising the first concerns that LLM performance may be affected by memorization of pre-training data. GPT-3 (Brown et al., 2020) scans data points having a 13-gram collision with anything in the pre-training Common Crawl (C4) dataset (Raffel et al., 2020). Some earlier benchmarks sourced from Wikipedia, the Children’s Book Test dataset (Hill et al., 2015), Quac (Choi et al., 2018) and SQuAD 2.0 (Rajpurkar et al., 2018), are nearly entirely contaminated. PaLM (Chowdhery et al., 2023) identifies 10 evaluation datasets at risk of contamination per their construction process, and for each dataset, partitions it into a clean and a contaminated subset based on whether at least 70% of the 8-grams of the data point can be found at least once in the training set. Both GPT-3 and PaLM authors conclude that the performance gap between clean sets and contaminated sets is mostly small across sets of various contamination degrees. GPT-4 (Achiam et al., 2023) measures cross-contamination between evaluation sets and pre-training data by computing 50-characters substring collisions. After examining GPT-4 performance gap between contaminated and clean subsets, the authors conclude that contamination overall has very little effect on the reported zero-shot results. Llama-2 (Touvron et al., 2023b) refines the methods above by using tokens instead of words to measure contamination, and asserts that a token is contaminated if it appears in any token n-gram longer than 10 tokens in both the evaluation sample and the training set. The formatting of prompts used for actual evaluation is also considered as it affects the performance more significantly. From there, the contamination level of a sample is defined as its percentage of contaminated tokens. The authors also introduce the usage of a skip-gram budget allowing matched spans between the evaluation sample and the training set to differ in at most four positions. Llama-3 (Dubey et al., 2024) follows a similar technique to Llama-2, but using 8-gram tokens instead of 10-gram. The ratio of contaminated tokens is varied to find the value with highest performance gain across models. Authors conclude that performance on PiQA (Bisk et al., 2020) and HellaSwag (Zellers et al., 2019) is affected by contamination. Recently, Qwen-2.5-Coder (Hui et al., 2024) removes all training data points with a 10-gram collision with the test set. Popular benchmarks HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), GSM8K (Cobbe et al., 2021), and MATH (Hendrycks et al., 2021) are affected.
Table 1 summarizes contaminated datasets found through all string-matching contamination detection described above. Notably, PIQA (Bisk et al., 2020), Winograd (Levesque et al., 2012), HumanEval (Chen et al., 2021a) and HellaSwag (Zellers et al., 2019) are flagged by at least two contamination detection techniques.
LLM String Matching Contaminated Datasets
GPT-2 (Radford et al., 2019) 8-gram words CoQA (Reddy et al., 2019), LAMBADA (Paperno et al., 2016)
GPT-3 (Brown et al., 2020) 13-gram words PIQA (Bisk et al., 2020), Winograd (Levesque et al., 2012), Children’s Book Test (Hill et al., 2015), WikiText-2 (Merity et al., 2016),
WikiText-103 (Merity et al., 2016), enwik8, text8
PaLM (Chowdhery et al., 2023) 8-gram words Winograd (Levesque et al., 2012), SQuAD-2 (Rajpurkar et al., 2018), WSC, ReCoRD, CB
GPT-4 (Achiam et al., 2023) 50-gram characters HumanEval (Chen et al., 2021a)
Llama-2 (Touvron et al., 2023b) 10-gram tokens HellaSwag (Zellers et al., 2019), MMLU (Hendrycks et al., 2020)
Llama-3 (Dubey et al., 2024) 8-gram tokens PiQA (Bisk et al., 2020), HellaSwag (Zellers et al., 2019), AGIEval (Zhong et al., 2023), BIG-Bench Hard (Srivastava et al., 2022),
BoolQ (Clark et al., 2019), OpenBookQA (Mihaylov et al., 2018), QuaC (Choi et al., 2018), SiQA (Sap et al., 2019)
Qwen-2.5-Coder (Hui et al., 2024) 10-gram words HumanEval (Chen et al., 2021a), MBPP (Austin et al., 2021), GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021)
Table 1: Datasets flagged with a high-level of contamination for several LLM pre-training studies, assessing contamination through string matching between the pre-training corpus and the evaluation samples.
Singh et al. (2024) review several string-matching contamination detection techniques, including the ones used by Brown et al. (2020) (Token-Match), Chowdhery et al. (2023) (Ngram-Match) and Touvron et al. (2023b) (Token-Extend). They introduce a new method called Longest-Match which measures the fraction of tokens part of the longest contaminated token span, avoiding the issue of frequent token spans artificially inflating contamination levels with other methods. Similarly to (Touvron et al., 2023b), Longest-Match also allows for skipgram. Longest-Match finds the highest Estimated Performance Gain (EPG) between contaminated and clean evaluation subsets. Through Llama-1 (Touvron et al., 2023a) and Pythia (Biderman et al., 2023) series of LLMs, the Llama pre-training corpus and The Pile (Gao et al., 2020) are found to be contaminated for many popular evaluation benchmarks, notably BigBench (Srivastava et al., 2022), HellaSwag (Zellers et al., 2019), PiQA (Bisk et al., 2020), MMLU (Hendrycks et al., 2020) and TriviaQA (Joshi et al., 2017).
2.1.2 Dataset-Level String Matching
@@
Other studies analyze overlap between the training split and the evaluation or test split of the same dataset to assess contamination levels. While standard practice in machine learning assumes that training, validation and test splits of the same dataset are disjoint sets sampled from the same distribution, in practice, for very large datasets it may happen that the training set still overlaps with the other splits. Dodge et al. (2021) investigate the C4 (Raffel et al., 2020) dataset and study benchmark data contamination, which measures how much training or test datasets from downstream NLP tasks appear in the C4 pre-training corpus. After studying both input-label and input contamination, they discover varied contamination, ranging from less than 2% to over 50%. Li (2023b) compute the METEOR (Banerjee & Lavie, 2005) score between matched pages from CommonCrawl and queries from Bing API and consider those with scores over 0.75 as contaminated. Detected contamination levels range from 1% on Winogrande (Sakaguchi et al., 2021) to 47% on C-Eval (Huang et al., 2023). Deng et al. (2023) propose to detect contamination by retrieving top-10 documents from pre-training datasets The Pile (Gao et al., 2020) or C4 (Raffel et al., 2020), then splitting retrieved documents into 13-gram chunks, and computing overlap metrics between these chunks and the evaluation data point’s chunks. TruthfulQA (Lin et al., 2021) exhibits high overlap with the pre-training datasets.
@@ 2.2 Embeddings Similarity
Lee et al. (2023) prevent contamination in their Open-Platypus dataset by removing test questions which have a cosine similarity (using SentenceTransformer embeddings (Reimers & Gurevych, 2019)) greater than 80% against any training item. Phi-1 (Gunasekar et al., 2023) runs a data contamination analysis between their CodeExercises dataset and the evaluation set HumanEval (Chen et al., 2021a). They show that embeddings-based retrieval between code snippets using L2 distance between normalized CodeGen-Mono 350M (Nijkamp et al., 2022) embeddings is effective, whereas n-gram overlap fails in the coding domain as it cannot capture the similarities in the logic between two coding functions.
The work from Riddell et al. (2024) provides a hybrid approach combining string matching and embeddings similarity: an aggregated score is designed as the maximum between the Levenshtein edit distance and the similarity from the Dolos toolkit (Maertens et al., 2022). Widely used code generation benchmarks MBPP (Austin et al., 2021) and HumanEval (Chen et al., 2021a) are contaminated with pre-training sets The Pile (Gao et al., 2020) and especially The Stack (Kocetkov et al., 2022), which significantly inflates performance of models like CodeGen-NL (Nijkamp et al., 2022).
2.3 Paraphrase Detection
Embeddings similarity is limited by its need to choose an appropriate similarity threshold. To improve robustness, some methods attempt to use LLMs themselves to detect contamination in LLMs, leveraging their high performance in zero-shot paraphrase detection (Witteveen & Andrews, 2019; Abaskohi et al., 2023).
Yang et al. (2023c) argue that contamination checks become challenging because of the presence of “rephrased samples”, which have the same semantics but different surface form as the original sample. Therefore, they first use embeddings similarity search to get the top-k similar samples with a given test sample, then prompt a strong LLM (namely, GPT-4) to examine whether any of the top-k samples is too close to the test case. Results show that this LLM Decontaminator algorithm works significantly better than existing methods at tagging contaminated samples.